# The Impact of Question Framing on the Performance of Automatic Occupation Coding

### v2

@@ 3.3 Analysis
H1. Question 1: Job Titles/Occupational Tasks Automatic Codability We use two readily available automatic coding tools for the German language, CASCOT and OccuCoDe, that can code answers into national German (KldB 2010) and international (ISCO-08) classifications. We code answers after the survey has been conducted, however OccuCoDe has a functionality to be embedded directly in the survey for coding during the survey. When either algorithm assigns a code to a job title or an occupational task, it produces a confidence score to indicate the level of certainty that a given occupation should receive a specific code. For example, OccuCoDe will assign an answer “surgeon” (Chirurg) to the German classification KldB 2010 code 81434 (Surgeon) with a confidence score of 0.8. The second most likely choice will be the code 81474 (Dentist) with a confidence score of 0.05 etc. CASCOT will assign the same job to the 81434 (Surgeon) with a confidence score of 70. The second most probable job code will be 81332 (Surgical technical assistant) with a confidence score 20, and so on. Both tools aim to offer high-quality suggestions and warn against relying on suggestions that have a low score. In practice, it means that if a person answers “Civil servant” (Beamter), OccuCoDe will suggest that the most likely job category is “Administrative assistant” (Verwaltungshelfer/in) with a confidence score of 0.46, CASCOT will offer “civil servant working at a stud farm” (Beamter/Beamtin Landgestüte) with the score 52. Both suggestions lie below the tools’ recommended minimum confidence score levels (0.535 for OccuCoDe (possible range: 0-1) (Simson et al., 2023) and 64 (possible range: 0-100) for CASCOT (Jones and Elias, 2004)). As a result, the job “civil servant” is considered not precise enough for automatic coding and will require a manual coder’s effort. Because each case that falls under the confidence cut-off is automatically flagged in the output file, the automatic tools are reliably error-prone: generic responses such as "civil servant" surface in a predictable way, allowing for targeted follow-up, a level of traceability rarely attainable with human coders. We will call occupations that can be confidently coded using automatic tools “easy-to-code” (CASCOT: confidence scores 64-100; OccuCoDe confidence scores 0.535-1); the remaining ones will be called “hard-to-code”(CASCOT: confidence scores 0-63; OccuCoDe confidence scores 0-0.5349) and include some non-codable occupations (score 0). Importantly, human coders will more often agree with easy-to-code answers and less often agree with hard-to code ones, meaning that the accuracy from hard-to-code answers is believed to be lower. We do not aim to compare the coding decisions of the two tools. Instead, we are only interested in whether more answers to the job title question (Question 1"Title") fall under the “easy-to-code” category than the ones asking about occupational task (Question 1"Task").
@@
Second, we preprocess the responses by tokenizing, removing stop words, and applying stemming. We then measure linguistic richness using the Type to Token Ratio (TTR), which evaluates the number of unique words relative to the total number of words. The TTR ranges from 0 to 1, with higher values indicating a greater proportion of unique words (deBoer, 2014; Gavras, 2022). We calculate separate TTR for answers to job titles and professional activities, measure the difference TTRdif = TTRprofessional activity - TTRjob titles, and apply bootstrapping (N=1000) to estimate whether the difference is statistically significant (= 0.05).
In addition, following Martínez et al. (2017); Struminskaya (2015); Meitinger et al. (2021), we compare job title responses with professional activity using the length of answers (in words and characters). We apply a two-sided Mann–Whitney U (MWU) test test to measure the difference in length, as well as Levene’s test to measure the variability in answer length.
@@
First, we compare the length of answers (in characters and words) to the Question 2 (detailed tasks) with the example and without the example (Martínez et al., 2017; Massing et al., 2019). We apply the one-sided MWU test and Levene’s test to measure the difference in answer length variation. Second, we used the same lexical categories defined in Hypothesis 2: nouns, nouns that represent an occupational title (ones that end with -mann, -frau, -er, -erin, -in, -eur, -e, -ist, -wirt), verbs, adjectives, stop words (German stop words like und, ich, der, die, das, etc.), derived nouns (nouns that are derived from verbs and have endings like -ung, -heit, -keit, -nis, -schaft). We again used Azure OpenAI to classify each response into these six categories.
@@ 4.1 Descriptive Statistics
We received 840 responses from the two agencies after excluding everyone who explicitly selected the option “I am retired, or I am not working.” As recorded in Table 1, two respondents did not answer a question (Hard non-response), and 63 respondents (7%) provided an answer that was classified as a Soft non-response (gibberish, irrelevant answer etc.) and also excluded from further analysis. All remaining responses were spellchecked and (where appropriate) brought to the same form (for example, Kfm Angestellster, kfm. Angestellter, kaufmännischer Angestellter –> kaufmännischer Angestellter) in order to exclude the effect of spelling on codability (Bao et al., 2020). The final number of observations with valid responses is 775. Respondents, on average, used 1 (Bilendi) to 1 (Forsa) words or from 15 to 16.7 characters to describe their job title (Quesions 1"Title") and from 1 (Bilendi) to 1 (Forsa) words or from 15 to 17 characters to answer the question about the occupational task (Question 1"Task").
@@ 4.2 Codability of answers
Hard-to-code responses included terms such as clerk, employee, self-employed, worker, consultant, and salesperson, among others. These are generic and ambiguous answers for which multiple occupational categories could be equally appropriate, as indicated by their low confidence scores. Both CASCOT and OccuCoDe showed a higher proportion of easy-to-code responses when participants answered the job title question. However, according to a series of chi-square (X2) tests, there was no statistically significant difference in the proportion of easy-to-code responses between Question 1"Title" and Question 1"Task". The 95% confidence intervals nonetheless indicate the range of effects still compatible with our data: for OccuCoDe (Forsa) the true difference could be anywhere from –3 to +19 percentage points, and for CASCOT the bands are –9 to +12 points in the Forsa sample and –8 to +10 points in Bilendi. Translating those limits into workload, switching to the job-title wording could, in the worst case (-9 pp), raise the manual-coding load by roughly 9% (about 12600 extra cases in a labour-force survey with 140000 interviews, the size of the German LFS). In the best-case (+19 pp) it could let about 26600 more answers be coded automatically. Since these bounds are wide, a larger study is needed before survey practitioners can choose the wording with firmer statistical confidence.
@@ 5 Conclusion
First, between 29% and 48% of occupational answers, depending on the question formulation, were confidently coded by the two automatic coding tools we tested. Second, both automatic tools performed better when coding responses to the "job title" question compared to the "occupational task" question. This was especially visible for OccuCoDe, where responses to the job title question had 8% higher codability rate. Third, contrary to our expectations, we found that irrespective of the question formulation, over 80% of respondents named their job title. However, respondents to the "occupational task" question more frequently provided generic or ambiguous titles that were not reliably codable by the two automatic tools. Fourth, in line with the findings of Martínez et al. (2017), the answers to the second question about detailed tasks and duties proved to be significantly longer if an example was provided. The response types also varied by question formulation, with answers to the example-based question containing more nouns and verbs (e.g., "I manage," "I lead"). However, examples steered respondents toward using certain words. The average vocabulary diversity per word did not differ between questions with and without examples.
@@
Moreover, we only tested two question wording bundles currently used in Germany - “berufliche Tätigkeit” ("occupational task") without examples, and “Berufsbezeichnung” ("job title") with examples. A larger, fully crossed design (including a hybrid "berufliche Tätigkeit" with examples and a more sparse version of "Berufsbezeichnung") could disentangle the distinct roles of stem wording and example provision. Future studies might also examine whether task-oriented examples (instead of current job-title examples) enhance coding performance for occupational task questions. Additionally, since our samples came from commercial panels, we cannot exclude the possibility that the very experienced respondents chose an easy path, providing minimal answers to open-ended occupational questions. These panels (treated equally in our analyses despite Forsa’s probability-based recruitment and Bilendi’s opt-in nature) may also underrepresent some hard-to-reach or specialized occupations. Future research should replicate this experiment in a non-panel study setting to address the above concerns.
@@
Our findings have several practical implications. The wording of the occupational question appears to impact data quality for manual and automatic coding, even if our evidence is preliminary. As noted by Velkoff et al. (2014), asking to name the job title might not be the best option for those with multiple employments or industry-specific jobs, but, according to our findings, the job title is the answer that will be most likely given. Researchers should pay attention to the wording of the occupational question when given the data for training and testing automatic occupational tools. Working with "job title -Berufsbezeichnung" instead of "occupational task -berufliche Tätigkeit" in German context may affect the accuracy of the automatic coding tools without adjustments to the algorithm. The findings might be different for other cultural and linguistic contexts; however, together with Martínez et al. (2017), our study points out that the origin of the data matters. Future studies should examine causal mechanism that trigger difference in responses.
@@
We note that our findings are most directly relevant to post-survey coding of open-ended responses. Following the recommendations from (Geis and Hoffmeyer-Zlotnik, 2000) and Hoffmeyer-Zlotnik et al. (2024), many German surveys ask for "occupational task" as a first question, forming a particular German tradition. By contrast, we recommend asking the "job title" question first. Our findings contradict a rationale from (Geis and Hoffmeyer-Zlotnik, 2000), not tested until now, who argued that the "occupational task" wording should yield better results. While we applied OccuCoDe for post-survey coding, the tool (like some other systems) can also be used interactively during data collection, allowing respondents to select from pre-classified categories. Such interactive approaches may change the coding dynamic (for example, respondents can revise their initial answer if the suggested categories are unsatisfactory), and represent an alternative option for researchers designing new surveys.
@@
Next, our study showed that working with only short job title types of answers is limiting. Following Christoph et al. (2020) that respondents don’t always know what is required of them to answer, we end up with many answers that are hard to code. Our results showed that detailed duties questions gave crucial information about some respondents’ managerial duties that could not be foreseen from their initial answers. Therefore, researchers in the field of automatic coding could aim not to reach the maximum level of detail with coding job titles but instead to expand allowed input to include detailed task descriptions or industry.
@@
Lastly, our study showed that asking respondents for detailed descriptions of their duties elicited longer responses. Previous research indicates that more extensive input benefited some automatic coding tools (Russ and et al., 2016; Nan Li and Bie, 2023). However, we also found that respondents follow the answer pattern and provide longer but similar answers using the same pool of words. Thus, researchers should not expect that their trained machine learning models are easily transferable and with equal performance from one survey to a different one because the respective question formulations in each survey will always influence what exactly respondents answer.
@@
Our study reconfirms what survey researchers have long known: question wording matters. Building consensus about how exactly researchers should ask for and code occupations in their respective surveys will need additional efforts. Especially as machine learning and AI methods spread, and coding tools are transferred between different surveys, it would be well worth it.
@@ 6 Acknowledgments
[ caption=Comparison of Census and Survey Questions in the USA and Germany (1950-2024), label=tab:appendix-1A, ] colspec=p1.5cm|p6cm|p6cm, rowhead=1 Years USA, ACS - American Community Survey P - population census [Ruggles et al., 2024] Germany, M - Microcensus P - population Census [Forschungsdatenzentrum, n.d., n.d., Volkszählung 1987, 1987, Zensus 2011, 2011, Zensus 2022, 2022]
@@
ACS: What was this person’s main occupation? Describe this person’s most important activities or duties. (after 2018) M: No change in question wording (until 2010)
@@
[ caption=Survey wordings of occupational task questions and whether examples are included. The table summarizes how various German surveys word their open-ended occupation items and whether examples or classification instructions are provided in the question stem or interviewer instructions, label=tab:appendix-1B, ] colspec=p2.5cm|p9.5cm|p1.5cm, rowhead=1 Survey / Instrument Exact wording Examples?
Demographic Standards (2024) [ADM Arbeitskreis Deutscher Markt- und Sozialforschungsinstitute e. V. et al., 2024] Welche berufliche Tätigkeit üben Sie aus? Wenn Sie nicht mehr Voll- oder Teilzeit erwerbstätig sind: Welche Tätigkeit haben Sie bei Ihrer früheren hauptberuflichen Erwerbstätigkeit zuletzt ausgeübt? No
Mikrozensus (2025) Welche Berufsbezeichnung hat Ihre gegenwärtige Tätigkeit?
Z. B.:
- Modeverkäufer/-in
- Grundschullehrer/-in
- Reiseverkehrskaufmann/-frau
- Bauingenieur/-in
- Elektronikmechaniker/-in
- Bauhilfsarbeiter/-in
- Krankenpfleger/-in Yes, in the stem of the question
ALLBUS 2023 [GESIS - Leibniz Institute for the Social Sciences, 2025] Welche berufliche Tätigkeit üben Sie in Ihrem Hauptberuf aus? Bitte beschreiben Sie Ihre berufliche Tätigkeit genau. No
NEPS 2024 [FDZ-LIfBi, 2024] Sagen Sie mir bitte, welche berufliche Tätigkeit Sie da ausgeübt haben! No
GLES 2025 Welche berufliche Tätigkeit übten Sie in Ihrem Hauptberuf aus? Hat diese Tätigkeit einen besonderen Namen?
@@
SOEP 2023 [INFAS, 2025] Welche berufliche Tätigkeit üben Sie derzeit aus?
@@
FReDA 2021 [INFAS, 2021] Welche berufliche Tätigkeit üben Sie derzeit hauptsächlich aus? Wenn Sie derzeit nicht erwerbstätig sind, welche berufliche Tätigkeit haben Sie zuletzt in Ihrer hauptsächlichen Erwerbstätigkeit ausgeübt?
Bitte geben Sie die genaue Tätigkeitsbezeichnung an, also z.B. nicht „kaufmännische/r Angestellte/r“, sondern „Speditionskaufmann bzw. -frau“, nicht „Arbeiter/in“, sondern „Maschinenschlosser/in“. Wenn Sie Beamter/in sind, geben Sie bitte Ihre Amtsbezeichnung an, z.B. „Polizeimeister/in“ oder „Studienrat/-rätin“. Wenn Sie Auszubildende/r sind, geben Sie bitte Ihren Ausbildungsberuf an. Yes, in the instructions
GEDA 2019/2020 [Robert Koch-Institut and Statistisches Bundesamt (Destatis)(2021), RKI] Welchen Beruf üben Sie derzeit hauptsächlich aus?
Geben Sie bitte die genaue Berufsbezeichnung an, nicht den Ausbildungsabschluss oder Rang. Zum Beispiel: - Blumenverkäuferin (nicht Verkäuferin) - Maurer (nicht Bauarbeiter) - Grundschullehrer (nicht Lehrer oder Beamte) - Unternehmensberaterin (nicht Betriebswirtin) Yes, in the instructions
GESIS Panel (from 2023 on) [GESIS - Leibniz Institute for the Social Sciences, 2023] Welchen Beruf üben Sie zur Zeit aus?
Bitte geben Sie die genaue Tätigkeitsbezeichnung an. Schreiben Sie z.B. "Speditionskauffrau" und nicht "kaufmännische Angestellte", oder "Maschinenschlosser" statt "Arbeiter". Wenn Sie Beamte/r sind, geben Sie bitte Ihre Amtsbezeichnung an, z.B. "Polizeimeister" oder "Studienrat". Wenn Sie Auszubildende/r sind, geben Sie bitte Ihren Ausbildungsberuf an. Yes, in the instructions

### v3

@@ 3.3 Analysis
H1. Question 1: Job Titles/Occupational Tasks Automatic Codability We use two readily available automatic coding tools for the German language, CASCOT and OccuCoDe, that can code answers into national German (KldB 2010) and international (ISCO-08) classifications. We code answers after the survey has been conducted, however OccuCoDe has a functionality to be embedded directly in the survey for coding during the survey. When either algorithm assigns a code to a job title or an occupational task, it produces a confidence score to indicate the level of certainty that a given occupation should receive a specific code. For example, OccuCoDe will assign an answer “surgeon” (Chirurg) to the German classification KldB 2010 code 81434 (Surgeon) with a confidence score of 0.8. The second most likely choice will be the code 81474 (Dentist) with a confidence score of 0.05 etc. CASCOT will assign the same job to the 81434 (Surgeon) with a confidence score of 70. The second most probable job code will be 81332 (Surgical technical assistant) with a confidence score 20, and so on. Both tools aim to offer high-quality suggestions and warn against relying on suggestions that have a low score. In practice, it means that if a person answers “Civil servant” (Beamter), OccuCoDe will suggest that the most likely job category is “Administrative assistant” (Verwaltungshelfer/in) with a confidence score of 0.46, CASCOT will offer “civil servant working at a stud farm” (Beamter/Beamtin Landgestüte) with the score 52. Both suggestions lie below the tools’ recommended minimum confidence score levels (0.535 for OccuCoDe (possible range: 0-1) (Simson2023) and 64 (possible range: 0-100) for CASCOT (jones2004cascot)). As a result, the job “civil servant” is considered not precise enough for automatic coding and will require a manual coder’s effort. Because each case that falls under the confidence cut-off is automatically flagged in the output file, the automatic tools are reliably error-prone: generic responses such as "civil servant" surface in a predictable way, allowing for targeted follow-up, a level of traceability rarely attainable with human coders. We will call occupations that can be confidently coded using automatic tools “easy-to-code” (CASCOT: confidence scores 64-100; OccuCoDe confidence scores 0.535-1); the remaining ones will be called “hard-to-code”(CASCOT: confidence scores 0-63; OccuCoDe confidence scores 0-0.5349) and include some non-codable occupations (score 0). Importantly, human coders will more often agree with easy-to-code answers and less often agree with hard-to code ones, meaning that the accuracy from hard-to-code answers is believed to be lower. We do not aim to compare the coding decisions of the two tools. Instead, we are only interested in whether more answers to the job title question (Question 1"Title") fall under the “easy-to-code” category than the ones asking about occupational task (Question 1"Task").
@@
Second, we preprocess the responses by tokenizing, removing stop words, and applying stemming. We then measure linguistic richness using the Type to Token Ratio (TTR), which evaluates the number of unique words relative to the total number of words. The TTR ranges from 0 to 1, with higher values indicating a greater proportion of unique words (deBoer2014; Gavs2022). We calculate separate TTR for answers to job titles and professional activities, measure the difference TTRdif = TTRprofessional activity - TTRjob titles, and apply bootstrapping (N=1000) to estimate whether the difference is statistically significant (= 0.05).
In addition, following Martinez2017; Struminskaya2015; meitinger2021, we compare job title responses with professional activity using the length of answers (in words and characters). We apply a two-sided Mann–Whitney U (MWU) test test to measure the difference in length, as well as Levene’s test to measure the variability in answer length.
@@
First, we compare the length of answers (in characters and words) to the Question 2 (detailed tasks) with the example and without the example (Martinez2017; Massing2019). We apply the one-sided MWU test and Levene’s test to measure the difference in answer length variation. Second, we used the same lexical categories defined in Hypothesis 2: nouns, nouns that represent an occupational title (ones that end with -mann, -frau, -er, -erin, -in, -eur, -e, -ist, -wirt), verbs, adjectives, stop words (German stop words like und, ich, der, die, das, etc.), derived nouns (nouns that are derived from verbs and have endings like -ung, -heit, -keit, -nis, -schaft). We again used Azure OpenAI to classify each response into these six categories.
@@ 4.1 Descriptive Statistics
We received 840 responses from the two agencies after excluding everyone who explicitly selected the option “I am retired, or I am not working.” As recorded in Table 1, two respondents did not answer a question (Hard non-response), and 63 respondents (7%) provided an answer that was classified as a Soft non-response (gibberish, irrelevant answer etc.) and also excluded from further analysis. All remaining responses were spellchecked and (where appropriate) brought to the same form (for example, Kfm Angestellster, kfm. Angestellter, kaufmännischer Angestellter –> kaufmännischer Angestellter) in order to exclude the effect of spelling on codability (bao2020occupation). The final number of observations with valid responses is 775. Respondents, on average, used 1 (Bilendi) to 1 (Forsa) words or from 15 to 16.7 characters to describe their job title (Quesions 1"Title") and from 1 (Bilendi) to 1 (Forsa) words or from 15 to 17 characters to answer the question about the occupational task (Question 1"Task").
@@ 4.2 Codability of answers
Hard-to-code responses included terms such as clerk, employee, self-employed, worker, consultant, and salesperson, among others. These are generic and ambiguous answers for which multiple occupational categories could be equally appropriate, as indicated by their low confidence scores. Both CASCOT and OccuCoDe showed a higher proportion of easy-to-code responses when participants answered the job title question. However, according to a series of chi-squared ($\chi^{2}$) tests, there was no statistically significant difference in the proportion of easy-to-code responses between Question 1"Title" and Question 1"Task". The 95% confidence intervals nonetheless indicate the range of effects still compatible with our data: for OccuCoDe (Forsa) the true difference could be anywhere from –3 to +19 percentage points, and for CASCOT the bands are –9 to +12 points in the Forsa sample and –8 to +10 points in Bilendi. Translating those limits into workload, switching to the job-title wording could, in the worst case (-9 pp), raise the manual-coding load by roughly 9% (about 12600 extra cases in a labour-force survey with 140000 interviews, the size of the German LFS). In the best-case (+19 pp) it could let about 26600 more answers be coded automatically. Since these bounds are wide, a larger study is needed before survey practitioners can choose the wording with firmer statistical confidence.
@@ 5 Conclusion
For more than a century, questionnaire designers have been thinking about the wording of occupational questions. It is now widely recommended to ask a sequence of two (or more) questions so that as many respondents as possible, even the hard-to-code ones, provide all the information needed for human coding. From this perspective, the goal is to maximize the information content of the complete sequence of questions. Questions at the start of the sequence developed under this perspective frequently steer respondents away from replying with their job titles. Another perspective emerges when considering automatic coding tools that have penetrated occupation coding in practice. Most existing tools rely on short inputs, particularly job titles, and may not reliably process longer, more detailed descriptions of tasks and duties. Since most respondents mention a job title in response to a starting question about their occupation, the information content from this first question is highly relevant for automatic coding approaches. We observe some possible tension between these two perspectives and, in practice, the need to decide whether the starting question should aim to collect job titles.
@@
First, between 30% and 48% of occupational answers, depending on the question formulation, were confidently coded by the two automatic coding tools we tested. Second, both automatic tools performed better when coding responses to the "job title" question compared to the "occupational task" question. This was especially visible for OccuCoDe, where responses to the job title question had 8% higher codability rate. Third, contrary to our expectations, we found that irrespective of the question formulation, over 80% of respondents named their job title. However, respondents to the "occupational task" question more frequently provided generic or ambiguous titles that were not reliably codable by the two automatic tools. Fourth, in line with the findings of Martinez2017, the answers to the second question about detailed tasks and duties proved to be significantly longer if an example was provided. The response types also varied by question formulation, with answers to the example-based question containing more nouns and verbs (e.g., "I manage," "I lead"). However, examples steered respondents toward using certain words. The average vocabulary diversity per word did not differ between questions with and without examples.
@@
Moreover, we only tested two question wording bundles currently recommended or used in Germany — “berufliche Tätigkeit” ("occupational task") without examples, and “Berufsbezeichnung” ("job title") with examples. A larger, fully crossed design (including a hybrid "berufliche Tätigkeit" with examples and a more sparse version of "Berufsbezeichnung") could disentangle the distinct roles of stem wording and example provision. Future studies might also examine whether task-oriented examples (instead of current job-title examples) enhance coding performance for occupational task questions. Future research should test the item-succession effects. For instance, asking about occupational tasks first might prime respondents to provide more specific job titles in a follow-up question. This hypothesis, implicit in the geis2000stand recommendation, remains untested and warrants future research. Additionally, since our samples came from commercial panels, we cannot exclude the possibility that the very experienced respondents chose an easy path, providing minimal answers to open-ended occupational questions. These panels (treated equally in our analyses despite Forsa’s probability-based recruitment and Bilendi’s opt-in nature) may also underrepresent some hard-to-reach or specialized occupations. Future research should replicate this experiment in a non-panel study setting to address the above concerns.
@@
Our findings have several practical implications. The wording of the occupational question appears to impact data quality for manual and automatic coding, even if our evidence is preliminary. As noted by ACS2014, asking to name the job title might not be the best option for those with multiple employments or industry-specific jobs, but, according to our findings, the job title is the answer that will be most likely given. Researchers should pay attention to the wording of the occupational question when given the data for training and testing automatic occupational tools. Working with "job title -Berufsbezeichnung" instead of "occupational task -berufliche Tätigkeit" in German context may affect the accuracy of the automatic coding tools without adjustments to the algorithm. The findings might be different for other cultural and linguistic contexts; however, together with Martinez2017, our study points out that the origin of the data matters. Future studies should examine causal mechanism that trigger difference in responses.
@@
We note that our findings are most directly relevant to post-survey coding of open-ended responses. Following the recommendations from geis2000stand and demographische2024standards, many German surveys ask for "occupational task" as a first question, forming a particular German tradition. Their occupational task question bundle is meant for coding into ISCO, but asking three open-ended questions in a survey is time-consuming and a challenge for any automatic coding solution. If there is only time to ask a single question or if the use of automatic coding is desired, our results suggest that the job title question will perform better than a single job task question, due to higher information density in job titles. While we applied OccuCoDe for post-survey coding, the tool (like some other systems) can also be used interactively during data collection, allowing respondents to select from pre-classified categories. Such interactive approaches may change the coding dynamic (for example, respondents can revise their initial answer if the suggested categories are unsatisfactory), and represent an alternative option for researchers designing new surveys.
@@
Next, our study showed that working with only short job title types of answers is limiting. Following Christoph2020, open-ended occupational questions do not always make sufficiently clear which aspects of occupational information are required, resulting in many answers that are hard to code. Our results showed that detailed duties questions gave crucial information about some respondents’ managerial duties that could not be foreseen from their initial answers. Therefore, researchers in the field of automatic coding could aim not to reach the maximum level of detail with coding job titles but instead to expand allowed input to include detailed task descriptions or industry.
@@
In addition, our study showed that asking respondents for detailed descriptions of their duties elicited longer responses. Previous research indicates that more extensive input benefited some automatic coding tools (Russ2016; li2023llm4jobs). However, we also found that respondents follow the answer pattern and provide longer but similar answers using the same pool of words. Thus, researchers should not expect that their trained machine learning models are easily transferable and with equal performance from one survey to a different one because the respective question formulations in each survey will always influence what exactly respondents answer.
Overall, our study experimentally explores possible variations of German occupation questions that geis2000stand and demographische2024standards recommend. Although we are not aware of any study that has fully implemented the lengthy three-question sequence proposed in these recommendations, they have been influential in the design of German occupation questions. Consequently, the wordings of the first occupation-related question in different surveys often resemble one another. However, we would like to highlight a number of differences and considerations: Firstly, different German surveys use the terms ‘Berufsbezeichnung’ (job title), ‘Beruf’ (occupation) and ‘berufliche Tätigkeit’ interchangeably. Secondly, many German surveys (e.g., NEPS, GLES, SOEP, FReDA) use the task-focused phrase ‘berufliche Tätigkeit’ from the recommendations. Their questions mention illustrative job titles in the instructions, clarifying how they expect respondents to answer. This is, however, contrary to the original intent of the phrase which was to steer respondents’ attention away from job titles and towards the tasks and duties performed in the job. Despite there being no such clarifying examples of job titles in our study, over 80% of our respondents mentioned a job title when asked about their ‘berufliche Tätigkeit’. This suggests that many respondents immediately want to answer with a job title when being asked about their occupation, regardless of the exact wording of the question. Thirdly, one line of reasoning expects higher quality from the job title question if a preceding item asks about detailed occupational tasks, while others reverse the question order, asking for job title first. Finally, if there is only time to ask a single question or if the use of automatic coding is desired, our results suggest that the job title question with examples will perform better than a single job task question without examples, due to a higher information density of job titles. While we applied OccuCoDe for post-survey coding, the tool (like some other systems) can also be used interactively during data collection, allowing respondents to select from pre-classified categories. These interactive approaches may alter the coding dynamic (for instance, respondents can revise their initial answer if the suggested categories are unsatisfactory), and offer an alternative option for researchers designing new surveys. We view these differences as a call to action for future research to explore which question formulation yields the best results.
@@
More generally, the results underscore what survey researchers have long known: question wording matters. Building greater consensus about how exactly researchers should ask for and code occupations in their respective surveys will need additional efforts. Especially as machine learning and AI methods spread, and coding tools are transferred between different surveys, it would be well worth it.
@@ 6 Acknowledgments
[ caption=Comparison of Census and Survey Questions in the USA and Germany (1950-2024), label=tab:appendix-1A, ] colspec=p1.5cm|p6cm|p6cm, rowhead=1 Years USA, ACS - American Community Survey P - population census (ipums_usa_v15) Germany, M - Microcensus P - population Census (forschungsdatenzentrum_mikrozensus; forschungsdatenzentrum_vz1970; zensus_1987; zensus2011; zensus2022)
@@
ACS: What was this person’s main occupation?
ACS: Describe this person’s most important activities or duties. (after 2018) M: No change in question wording (until 2010)
@@
[ caption=Survey wordings of occupational task questions and whether examples are included. The table summarizes how various German surveys word their open-ended occupation items and whether examples or classification instructions are provided in the question stem or interviewer instructions, label=tab:appendix-1B, ] colspec=p2.5cm|p4.5cm|p4.5cm|p1.5cm, rowhead=1 Survey / Instrument First question wording Second or third question wording (if applicable) Examples?
Demographic Standards (2024) (demographische2024standards) Welche berufliche Tätigkeit üben Sie aus? Wenn Sie nicht mehr Voll- oder Teilzeit erwerbstätig sind: Welche Tätigkeit haben Sie bei Ihrer früheren hauptberuflichen Erwerbstätigkeit zuletzt ausgeübt? Second question: Bitte beschreiben Sie diese berufliche Tätigkeit genau.
Third question: Hat dieser Beruf noch einen besonderen Namen? 1: ja und zwar: / nein No
Mikrozensus (online version) (2023) (bernhard_hochstetter_berufsklassifizierung_nodate) Welche Berufsbezeichnung hat Ihre gegenwärtige Tätigkeit? Z. B.: Modeverkäufer/-in; Grundschullehrer/-in; Reiseverkehrskaufmann/-frau; Bauingenieur/-in; Elektronikmechaniker/-in; Bauhilfsarbeiter/-in; Krankenpfleger/-in Yes, in the stem of the question
Mikrozensus (paper version) (2025) (mikro2025) Bitte beschreiben Sie Ihre gegenwärtige Tätigkeit in Stichworten. z.B. Verkauf von Kleidung; Kinder an der Grundschule unterrichten; Kundinnen und Kunden über Reiseangebote beraten und informieren; Bauwerke im Hochbau entwerfen oder planen; Elektronische Schaltungen aufbauen und prüfen; Beton, Gips und Mörtel mischen; Patientinnen und Patienten (vor, während und nach Operationen) betreuen und versorgen
Second question: Welche Berufsbezeichnung hat Ihre gegenwärtige Tätigkeit? z.B. Modeverkäufer/-in; Grundschullehrer/-in; Reiseverkehrskaufmann/-frau; Bauingenieur/-in; Elektronikmechaniker/-in; Bauhilfsarbeiter/-in; Krankenpfleger/-in Yes, in the stem of the question
ALLBUS 2023 (GESIS_ALLBUS_2023_CAWI) Welche berufliche Tätigkeit üben Sie in Ihrem Hauptberuf aus? Bitte beschreiben Sie Ihre berufliche Tätigkeit genau. Hat dieser Beruf, diese Tätigkeit noch einen besonderen Namen? No
NEPS 2024 (FDZLIfBi_NEPS_SC6_W15_2024) Sagen Sie mir bitte, welche berufliche Tätigkeit Sie da ausgeübt haben! According to interviewer instructions an exact job title is required. Können Sie mir diese berufliche Tätigkeit noch konkreter benennen? Hat das, was Sie gemacht haben bzw. machen, noch eine genauere Bezeichnung? No
GLES 2025 (GESIS_GLES_RCS_2025) Welche berufliche Tätigkeit übten Sie in Ihrem Hauptberuf aus? Hat diese Tätigkeit einen besonderen Namen?
@@
SOEP 2023 (SOEP2025) Welche berufliche Tätigkeit üben Sie derzeit aus?
@@
FReDA 2021 (FReDA_W1B_2021) Welche berufliche Tätigkeit üben Sie derzeit hauptsächlich aus? Wenn Sie derzeit nicht erwerbstätig sind, welche berufliche Tätigkeit haben Sie zuletzt in Ihrer hauptsächlichen Erwerbstätigkeit ausgeübt?
Bitte geben Sie die genaue Tätigkeitsbezeichnung an, also z.B. nicht „kaufmännische/r Angestellte/r“, sondern „Speditionskaufmann bzw. -frau“, nicht „Arbeiter/in“, sondern „Maschinenschlosser/in“. Wenn Sie Beamter/in sind, geben Sie bitte Ihre Amtsbezeichnung an, z.B. „Polizeimeister/in“ oder „Studienrat/-rätin“. Wenn Sie Auszubildende/r sind, geben Sie bitte Ihren Ausbildungsberuf an. Hat dieser Beruf noch eine besondere Bezeichnung? Yes, in the instructions
GEDA 2019/2020 (RKI_GEDA_2019_2020_EHIS) Welchen Beruf üben Sie derzeit hauptsächlich aus?
Geben Sie bitte die genaue Berufsbezeichnung an, nicht den Ausbildungsabschluss oder Rang. Zum Beispiel: - Blumenverkäuferin (nicht Verkäuferin) - Maurer (nicht Bauarbeiter) - Grundschullehrer (nicht Lehrer oder Beamte) - Unternehmensberaterin (nicht Betriebswirtin) Um die Einordnung Ihres Berufes zu erleichtern, geben Sie bitte zusätzliche Erläuterungen in Stichworten an. Zum Beispiel: als Blumenverkäuferin: Kundenberatung, Verkauf, Verpacken von Pflanzen; als Zollbeamter: Zollfahndung, Einsatzplanung, Pressearbeit; Falls Sie Führungsaufgaben wahrnehmen, vermerken Sie dies auch.; als KFZ-Mechaniker: Wartung, Instandsetzung, Ausrüstung von Kraftfahrzeugen, Leitung der Werkstatt Yes, in the instructions
GESIS Panel (from 2023 on) (GESIS_Panel_ZG_2023) Welchen Beruf üben Sie zur Zeit aus?
Bitte geben Sie die genaue Tätigkeitsbezeichnung an. Schreiben Sie z.B. "Speditionskauffrau" und nicht "kaufmännische Angestellte", oder "Maschinenschlosser" statt "Arbeiter". Wenn Sie Beamte/r sind, geben Sie bitte Ihre Amtsbezeichnung an, z.B. "Polizeimeister" oder "Studienrat". Wenn Sie Auszubildende/r sind, geben Sie bitte Ihren Ausbildungsberuf an. Bitte beschreiben Sie noch kurz lhre berufliche Tätigkeit und Ihre wesentlichen Aufgaben. Beispiele sind "Regale mit Produkten auffüllen und Inventur machen", "Patienten versorgen, Medikamente geben, Vitalzeichen überwachen", "Patienten betreuen, Zähne und Zahnfleisch behandeln", "Überwachung des Warenbestands in der Damenabteilung, Bedienen von Kundinnen und Kassiertätigkeit". Yes, in the instructions